At sea level, a scuba diver’s tank can hold 100 liters of air at standard pressure and room temperature (20 °C). Once he travels 30 meters deep, the volume decreases 25% to 75 liters and the water temperature falls to standard temperature. What is the pressure of the air in the diver’s cylinder when he is 30 meters deep, assuming that the mass remains constant?
1
Expert's answer
2019-10-09T07:37:06-0400

Solution.

To determine the air pressure in the driver cylinder at a depth of 30 m, we write down the klayperon-Mendleev equation:

"\frac{p1 \times V1}{T1} = \frac{p2 \times V2}{T2}"

"p2 = \frac{p1 \times V1 \times T2}{V2 \times T1}"

"p2 = \frac{101325 \times 100 \times 273}{75 \times 293} = 125878.16 \ Pa"

Answer:

p2 = 125878.16 Pa